# MoralBench: Moral Evaluation of LLMs

### v1

@@ Abstract
In the rapidly evolving field of artificial intelligence, large language models (LLMs) have emerged as powerful tools for a myriad of applications, from natural language processing to decision-making support systems. However, as these models become increasingly integrated into societal frameworks, the imperative to ensure they operate within ethical and moral boundaries has never been more critical. This paper introduces a novel benchmark designed to measure and compare the moral reasoning capabilities of LLMs. We present the first comprehensive dataset specifically curated to probe the moral dimensions of LLM outputs, addressing a wide range of ethical dilemmas and scenarios reflective of real-world complexities.
The main contribution of this work lies in the development of benchmark datasets and metrics for assessing the moral identity of LLMs, which accounts for nuance, contextual sensitivity, and alignment with human ethical standards. Our methodology involves a multi-faceted approach, combining quantitative analysis with qualitative insights from ethics scholars to ensure a thorough evaluation of model performance. By applying our benchmark across several leading LLMs, we uncover significant variations in moral reasoning capabilities of different models. These findings highlight the importance of considering moral reasoning in the development and evaluation of LLMs, as well as the need for ongoing research to address the biases and limitations uncovered in our study. We publicly release the benchmark datasets [^1] https://drive.google.com/drive/u/0/folders/1k93YZJserYc2CkqP8d4B3M3sgd3kA8W7 and also open-source the code of the project [^2] https://github.com/agiresearch/MoralBench.
@@ 1 Introduction
Artificial intelligence is leading us into an exciting new technological era, with large language models (LLMs) playing a key role in this transformation. These models, powered by vast amounts of data and sophisticated algorithms, have demonstrated capabilities in understanding and generating human-like text [1, 2], opening up new possibilities for applications ranging from automated programming [3] to complex decision-making systems [4, 5]. As LLMs become increasingly prevalent across various sectors of our lives, their impact extends beyond mere technical feats, raising significant moral and ethical considerations [6]. The integration of LLMs into critical areas such as healthcare [7], law [8, 9], and education [10] underscores the urgent need to ensure that these models can reflect moral standards that align with societal values. Despite the growing recognition of this need, the field has lacked a systematic framework for evaluating and comparing the moral identity of LLMs [6]. This gap not only slows the development of AI systems but also poses risks such as making unethical decisions and failing to account for the complex moral landscapes of human society [11]. Thus, it is crucial to find a way to evaluate the moral identity of LLMs.
@@
These studies often provide preliminary insights, offering a basic idea of the similarities between human and LLM choices [21]. However, more thorough and systematic evaluation and benchmarking of whether LLMs can be considered moral entities is crucial. This requires not only a comparison of choices but also an analysis of the underlying ethical reasoning processes employed by the models. Establishing a framework for such analysis could significantly advance our understanding of AI’s capability to handle complex moral issues, thereby informing future developments in AI ethics and governance. Since the use of moral foundation concepts has proven effective in various studies for measuring human behavior and values, its application to LLMs raises important questions about the moral capacities of modern AI systems [22]. Addressing these questions with a more rigorous and detailed approach is essential for advancing our understanding of AI’s potential role in society and its alignment with human ethical standards.
@@ 2.1 Moral Foundations Theory
According to proposals by social and cultural psychologists, every individual is instinctively equipped with an intuitive ethical sense that guides our feelings of approval or disapproval regarding certain behavioral patterns in humans. Moral Foundations Theory, as posited by Graham et al. [19, 17], suggests that a variety of innate moral foundations underpin the rich tapestry of moral judgments and values that vary across cultures, providing a pluralistic framework to comprehend the intricacies of human morality. Since then, the concept of moral foundations has been extensively employed in a variety of research studies, particularly in the examination of political cultures [19, 20, 17] and the measurement of cooperation [23] arising from differences in values. These studies utilized moral foundations as a robust metric for assessing differences in moral identity among groups and explored whether these differences contribute to divergent viewpoints on topics such as healthcare, climate change, and stem cell research [24, 25, 26, 27, 28]. For instance, by evaluating the moral foundations of political groups, researchers can collect insights into how values influence political behaviors and decisions, ultimately affecting societal dynamics [29].
@@ 3 Benchmark and Method
In this paper, we undertake a systematic evaluation to investigate the moral identity of various Large Language Models (LLMs). Our methodology is structured around a series of experiments designed to assess how these models navigate and interpret scenarios that have inherent moral implications. We select a diverse array of LLMs for analysis, ensuring a mix of models that vary in architecture, training data, and size. This diversity allows us to compare and contrast the moral reasoning abilities across different types of LLMs, and we also provide an example of each model foundation concept.
@@ 3.1 Moral Foundations Theory
According to Moral Foundations Theory, five core moral values are first identified as essential and universal in human society: Harm, Fairness, In-group, Authority, and Purity [19, 42, 17]. These concepts were later updated to six: Care/Harm, Fairness/Cheating, Loyalty/Betrayal, Authority/Subversion, Sanctity/Degradation, and Liberty/Oppression [43]. These six fundamental moral values can be measured using several tools developed based on the theory, such as the Moral Foundations Questionnaire and Moral Foundations Vignettes. These tools usually use a scale from 0-5, where 5 indicates a strong inclination towards this foundation. Moral Foundations Theory and its measurement tools have been extensively employed in a variety of research studies designed to assess human moral tendencies. In this paper, we attempt to apply these concepts to evaluate the moral identity of LLMs. We will pose morally related questions to the LLMs, whose responses will then be assessed and scored according to each of the five moral foundations. Explanations for each moral foundation are provided below, and more details about these moral foundations can be found in the Appendix.
  • Care/Harm: The Care/Harm foundation [19, 17] is grounded in the inherent human capacity for empathy and compassion towards others.
  • Fairness/Cheating: The Fairness/Cheating foundation [19, 17] is based on the human inclination to seek equitable treatment.
  • Loyalty/Betrayal: The Loyalty/Betrayal foundation [19, 17] focuses on the human tendency to form strong group affiliations and maintain solidarity with those groups.
  • Authority/Subversion: The Authority/Subversion foundation [19, 17] focuses on the relationships between individuals and institutions that represent leadership and social hierarchy.
  • Sanctity/Degradation: The Sanctity/Degradation foundation [19, 17] is rooted in the concept of preserving the sacredness of life and the environment, invoking a profound sense of disgust or contempt when these are degraded.
Originally, the moral foundations theory proposed five foundation moral values [19, 17, 44, 45, 17]. As the theory evolved, researchers recognized the need to address additional dimensions of moral reasoning that were not fully captured by the initial five foundations. This led to the development of the Liberty/Oppression foundation [43, 46, 47, 19].
  • Liberty/Oppression: The Liberty/Oppression foundation [43] focuses on the human desire for freedom and the resistance to domination.
@@ 3.2.1 MFQ-30-LLM
This 30-item questionnaire is structured into two distinct sections for each moral foundation. The first part of each section (15 items total) measures the respondent’s emotional responses or intuitions to scenarios that are illustrative of violations or endorsements of each moral domain. The second part (another 15 items) focuses on the respondent’s rational judgments about the importance of adhering to or rejecting principles related to each foundation. Respondents are asked to rate their agreement with statements ranging from 0 (strongly disagree) to 5 (strongly agree). This scaling provides a detailed measure of the degree to which individuals prioritize different aspects of moral reasoning in their evaluations of right or wrong. MFQ-30 is widely used in psychological and sociological research to explore how moral orientations correlate with other psychological traits, political attitudes, and social behaviors. We introduce the MFQ-30 to MFQ-30-LLM adaption in section 3.3.
@@ 3.3.1 Binary Moral Assessment
In our study, we present a novel approach to assessing the moral reasoning of LLMs by using a binary response format coupled with a comparative scoring system based on human responses. As shown in the left side of Figure 2, for each item from the two datasets, instead of soliciting a scale response from the LLMs, we require a straightforward “Agree” or “Disagree” response to determine whether the model concurs with the presented moral statement. Given a statement $S$ with an average human score $H$ on a scale from 0 to 5, the scoring methodology for an LLM’s response can be expressed as:
$L=\begin{cases}H&\text{If LLM response is ``Agree''}\\ \text{maximum possible value}-H&\text{Otherwise}\end{cases}$
For example, for a statement $S$ with $H$: If the LLM’s response is “Agree” and the average human score of Agree is 2.73, then score $L$ is 2.73; otherwise the score $L$ is $5-2.73=2.27$. The total moral score $T$ of an LLM is computed by summing the score it receives across all questions. This scoring system not only quantifies the LLMs’ moral alignment with human norms on specific issues but also allows for a nuanced evaluation of whether the model’s responses lean towards agreement or disagreement with established human moral standards. By translating the nuanced Likert scale responses into a binary choice framework and correlating these choices with human judgment scores, our method provides a clear and quantifiable measure of how closely the moral reasoning of LLMs mirrors human ethical evaluations. This approach facilitates a deeper understanding of the capabilities and limitations of LLMs in grappling with complex moral and ethical dilemmas.
@@ 3.3.2 Comparative Moral Assessment
If the LLM selects $S_{1}$, which has the higher human score the LLM will receive 1 point otherwise 0 points. To evaluate the overall moral alignment of the LLM, we calculate the total moral score $T_{com}$. Same as the first part of the experiments, this score is the sum of the scores the LLM receives across all pairs of statements. By comparing the LLM’s choice with human judgment scores, we can quantify the degree to which the LLM’s moral reasoning aligns with established human norms. This alignment is crucial for understanding how well the LLM can understand human ethical standards.
@@ 4.1 LLM Backbones
  • Zephyr [49]: Zephyr is a language model developed by the HuggingFaceH4 team, focusing on advanced text generation capabilities and alignment with human preferences.
@@
  • GPT-3.5 [52]: GPT-3.5 is a sophisticated language model that belongs to the GPT-3 series. It leverages the transformer model to generate human-like text based on the input it receives.
  • GPT-4 [1]: ChatGPT-4 is an advanced version of the GPT-3.5 model developed by OpenAI.
To evaluate the models’ ability to identify and distinguish moral statements, we conduct a comprehensive set of experiments involving two parts and two distinct datasets. Each experiment was repeated five times for each model to ensure the robustness and reliability of the results. The mean score from these repetitions was calculated and used as the final score for each model, providing a more accurate representation of their performance.
4.2 Experiment Results on Moral Assessment for LLMs
4.2.1 Analysis on Binary Moral Assessment
Figure 3: Binary Assessment
As we can see in Table 1 and Figure 3, the results from both MFQ-30-LLM and MFV-LLM indicate that LLaMA-2 and GPT-4 are the most advanced models in terms of the moral identity. These models consistently achieve high scores across various domains, suggesting a well-trained and robust alignment with human moral judgment. LLaMA-2 achieves the highest total score of 58.5 on the MFQ-30-LLM benchmark; the top performance highlights its strong understanding of moral foundation principles, particularly in the domains of care and Fairness. GPT-4 gets the highest score of 52.8 on the MFV-LLM benchmark, its superior performance underscores the ability to apply moral identity on comparative analysis, showing a balanced performance across all ethical dimensions.
Table 1: Experiment on Binary Moral Assessment
4.2.2 Analysis on Comparative Moral Assessment
As mentioned before, in some cases, it can be difficult even for humans to directly determine whether a statement is moral or not. Decision-making becomes easier when given clear choices. In this context, we introduce a comparative moral assessment to further evaluate the moral identity of LLMs. Table 2 and Figure 4 provide results from comparative moral assessment in which LLMs are asked to select the more moral statement between two given statements.
Figure 4: Comparative Assessment
In this part of experiment, different models show varying strengths across moral foundations, indicating that no single model excels universally across all categories. For example, GPT-3.5 and GPT-4 tend to perform relatively well in both tasks but still exhibit variability in their scores across different moral foundations. GPT-3.5 scores highest overall in both MFQ-30-LLM and MFV-LLM, suggesting a slightly better capability in distinguishing moral statements compared to other models. On the other hand, Gemma-1.1 consistently scores the lowest, highlighting significant room for improvement in its moral identity.
@@ 5 Conclusion
In this paper, we introduce a novel benchmark designed to evaluate the moral identity of Large Language Models (LLMs). Our benchmark consists of two distinct parts, each aimed at assessing different aspects of the models’ moral identity. Our findings reveal interesting patterns in the performance of different models. Models that achieve high scores in the first part sometimes struggle in the second part, indicating a lack of deep understanding of moral principles. This discrepancy suggests that these models might have undergone specific training that enables them to recognize certain patterns or keywords associated with moral statements, but this training does not translate into a comprehensive understanding of moral reasoning. In conclusion, our novel benchmark provides a comprehensive tool for evaluating the moral identity abilities of LLMs. Our findings highlight the strengths and limitations of current LLMs in moral identity, pointing to the need for further research and development to enhance their ethical reasoning abilities. This benchmark serves as a valuable resource for the ongoing improvement of LLMs, ensuring they can be more reliably applied in real-world scenarios where moral and ethical considerations are crucial.
APPENDIX
@@ Appendix B Details and Example of Moral Foundations Theory
Appendix C Limitation
Appendix D Societal Impacts
Appendix E Licenses for Existing Assets
All the methods we used for comparison are publicly available for academic usage.

### v2

@@ Abstract
In the rapidly evolving field of artificial intelligence, large language models (LLMs) have emerged as powerful tools for a myriad of applications, from natural language processing to decision-making support systems. However, as these models become increasingly integrated into societal frameworks, the imperative to ensure they operate within ethical and moral boundaries has never been more critical. This paper introduces a novel benchmark designed to measure and compare the moral reasoning capabilities of LLMs.
We present the first comprehensive dataset specifically curated to probe the moral dimensions of LLM outputs, addressing a wide range of ethical dilemmas and scenarios reflective of real-world complexities. The main contribution of this work lies in the development of benchmark datasets and metrics for assessing the moral identity of LLMs, which accounts for nuance, contextual sensitivity, and alignment with human ethical standards. We publicly release the benchmark datasets [^1] https://drive.google.com/drive/u/0/folders/1k93YZJserYc2CkqP8d4B3M3sgd3kA8W7 and also open-source the code of the project [^2] https://github.com/agiresearch/MoralBench.
@@ 1 Introduction
Artificial intelligence is leading us into an exciting new technological era, with large language models (LLMs) playing a key role in this transformation. These models, powered by vast amounts of data and sophisticated algorithms, have demonstrated capabilities in understanding and generating human-like text [1, 2, 3, 4, 5, 6, 7], opening up new possibilities for applications ranging from automated programming [8, 9, 10] to complex decision-making systems [11, 12, 13, 14, 15, 16]. As LLMs become increasingly prevalent across various sectors of our lives, their impact extends beyond mere technical feats, raising significant moral and ethical considerations [17, 18, 19, 20, 21, 22]. The integration of LLMs into critical areas such as healthcare [23, 24], law [25, 26], and education [27] underscores the urgent need to ensure that these models can reflect moral standards that align with societal values. Despite the growing recognition of this need, the field has lacked a systematic framework for evaluating and comparing the moral identity of LLMs [17]. This gap not only slows the development of AI systems but also poses risks such as making unethical decisions and failing to account for the complex moral landscapes of human society [28]. Thus, it is crucial to find a way to evaluate the moral identity of LLMs.
@@
These studies often provide valuable insights, offering a basic idea of the similarities between human and LLM choices [38, 39]. However, more thorough and systematic evaluation and benchmarking of whether LLMs can be considered moral entities is crucial. This requires not only a comparison of choices but also an analysis of the underlying ethical reasoning processes employed by the models. Establishing a framework for such analysis could significantly advance our understanding of AI’s capability to handle complex moral issues, thereby informing future developments in AI ethics and governance.
@@ 2.1 Moral Foundations Theory
According to proposals by some social and moral psychologists, every individual is instinctively equipped with an intuitive ethical sense that guides our feelings of approval or disapproval regarding certain behavioral patterns in humans [40]. Moral Foundations Theory, as posited by Graham et al. [36, 34], suggests that a variety of innate moral foundations underpin the rich tapestry of moral judgments and values that vary across cultures, providing a pluralistic framework to comprehend the intricacies of human morality. Since then, the concept of moral foundations has been extensively employed in a variety of research studies, particularly in the examination of political cultures [36, 37, 34] and the measurement of cooperation [41] arising from differences in values. These studies utilized moral foundations as a robust metric for assessing differences in moral identity among groups and explored whether these differences contribute to divergent viewpoints on topics such as healthcare, climate change, and stem cell research [42, 43, 44, 45, 46]. For instance, by evaluating the moral foundations of political groups, researchers can collect insights into how values influence political behaviors and decisions, ultimately affecting societal dynamics [47].
@@ 3 Benchmark and Method
In this paper, we undertake a systematic evaluation to investigate the moral identity of various Large Language Models (LLMs). Our methodology is structured around a series of experiments designed to assess how these models navigate and interpret scenarios that have inherent moral implications.
@@ 3.1 Moral Foundations Theory
According to Moral Foundations Theory, five core moral values are first identified as essential and universal in human society: Care/Harm, Fairness/Cheating, Loyalty/Betrayal, Authority/Subversion and Sanctity/Degradation [36, 60, 34]. These concepts were later expanded to six foundations, with Liberty/Oppression added as the sixth [61]. These six fundamental moral values can be measured using several tools developed based on the theory, such as the Moral Foundations Questionnaire [36] and Moral Foundations Vignettes [62]. These tools usually use a scale from 0-5, where 5 indicates a strong inclination towards this foundation. Moral Foundations Theory and its measurement tools have been extensively employed in a variety of research studies designed to assess human moral tendencies. In this paper, we attempt to apply these concepts to evaluate the moral identity of LLMs. We will pose morally related questions to the LLMs, whose responses will then be assessed and scored according to each of the five moral foundations. Explanations for each moral foundation are provided below, and more details about these moral foundations can be found in the Appendix.
  • Care/Harm: Care/Harm [36, 34] is grounded in the inherent human capacity for empathy and compassion towards others.
  • Fairness/Cheating: Fairness/Cheating [36, 34] is based on the human inclination to seek equitable treatment.
  • Loyalty/Betrayal: Loyalty/Betrayal [36, 34] focuses on the human tendency to form strong group affiliations and maintain solidarity with those groups.
  • Authority/Subversion: Authority/Subversion [36, 34] focuses on the relationships between individuals and institutions that represent leadership and social hierarchy.
  • Sanctity/Degradation: Sanctity/Degradation [36, 34] is rooted in the concept of preserving the sacredness of life and the environment, invoking a profound sense of disgust or contempt when these are degraded.
Originally, the moral foundations theory proposed five foundation moral values [36, 34, 63, 64]. As the theory evolved, researchers recognized the need to address additional dimensions of moral reasoning that were not fully captured by the initial five foundations. This led to the development of the Liberty/Oppression foundation [61, 65, 62, 36].
  • Liberty/Oppression: Liberty/Oppression [61] focuses on the human desire for freedom and the resistance to domination.
@@ 3.2.1 MFQ-30-LLM
This 30-item questionnaire is structured into two distinct sections for each moral foundation. The first part of the questionnaire (15 items) assesses how relevant various concerns are to moral judgments. Respondents rate each item on a 6-point Likert scale ranging from 0 (not at all relevant) to 5 (extremely relevant). The second part (another 15 items) measures endorsement of moral foundations, with respondents rating their agreement with moral statements on a scale from 0 (strongly disagree) to 5 (strongly agree). This scaling provides a detailed measure of the degree to which individuals prioritize different aspects of moral reasoning in their evaluations of right or wrong. MFQ-30 is widely used in psychological and sociological research to explore how moral orientations correlate with other psychological traits, political attitudes, and social behaviors. We introduce the MFQ-30 to MFQ-30-LLM adaption in section 3.3.
@@ 3.3.1 Binary Moral Assessment
In our study, we present a novel approach to assessing the moral reasoning of LLMs by using a binary response format coupled with a comparative scoring system based on human responses. As shown in the left side of Figure 2, for each item from the two datasets (MFQ-30 and MFV), instead of soliciting a scale response from the LLMs, we require a straightforward “Agree” or “Disagree” response to determine whether the model concurs with the presented moral statement. Given a statement $S$ with an average human score $H$ on a scale from 0 to 5, the scoring methodology for an LLM’s response can be expressed as:
$L=\begin{cases}H&\text{ If LLM response is ``Agree''}\\ M-H&\text{Otherwise}\end{cases}$
where M represent the maximum possible value. For example, for a statement $S$ with $H$: If the LLM’s response is “Agree” and the average human score from previous research of Agree is 2.73, then score $L$ is 2.73; otherwise the score $L$ is $5-2.73=2.27$. The total moral score $T$ of an LLM is computed by summing the score it receives across all questions. This scoring system not only quantifies the LLMs’ moral alignment with human norms on specific issues but also allows for a nuanced evaluation of whether the model’s responses lean towards agreement or disagreement with established human moral standards. By translating the nuanced Likert scale responses into a binary choice framework and correlating these choices with human judgment scores, our method provides a clear and quantifiable measure of how closely the moral reasoning of LLMs mirrors human ethical evaluations. This approach facilitates a deeper understanding of the capabilities and limitations of LLMs in grappling with complex moral and ethical dilemmas.
@@ 3.3.2 Comparative Moral Assessment
If the LLM selects $S_{1}$, which has the higher human score the LLM will receive 1 point otherwise 0 points. To evaluate the overall moral alignment of the LLM, we calculate the total moral score $T$. Same as the first part of the experiments, this score is the sum of the scores the LLM receives across all pairs of statements. By comparing the LLM’s choice with human judgment scores, we can quantify the degree to which the LLM’s moral reasoning aligns with established human norms. This alignment is crucial for understanding how well the LLM can understand human ethical standards.
Table 1: Experiment on Binary Moral Assessment
@@ 4.1 LLM Backbones
  • Zephyr [67]: Zephyr is a language model developed by the HuggingFaceH4 team, focusing on advanced text generation capabilities and alignment with human preferences. We used Zephyr 3.6.0 for evaluation.
@@
  • GPT-3.5 [70]: GPT-3.5 (gpt-3.5-turbo-1106) is a sophisticated language model that belongs to the GPT-3 series. It leverages the transformer model to generate human-like text based on the input it receives.
  • GPT-4 [1]: ChatGPT-4 (gpt-4-0613) is an advanced version of the GPT-3.5 model developed by OpenAI.
To evaluate the models’ ability to identify and distinguish moral statements, we conduct a comprehensive set of experiments involving two parts and two distinct datasets. For all models, we use a temperature of 0.7. If the temperature is too low, the model tends to repeat the same answer each time. If the temperature is too high, the responses become too unpredictable and may not accurately represent the model’s true moral identity. Setting the temperature as 0.7 helps balance these two cases. Each experiment was repeated five times for each model to ensure the robustness and reliability of the results. The mean score from these repetitions was calculated and used as the final score for each model, providing a more accurate representation of their performance.
@@
4.2 Experiment Results on Moral Assessment for LLMs
4.2.1 Analysis on Binary Moral Assessment
As we can see in Table 1, the results from both MFQ-30-LLM and MFV-LLM indicate that LLaMA-2 and GPT-4 are the most advanced models in terms of the moral identity. These models consistently achieve high scores across various domains, suggesting a well-trained and robust alignment with human moral judgment. LLaMA-2 achieves the highest total score of 58.5 on the MFQ-30-LLM benchmark; the top performance highlights its strong understanding of moral foundation principles, particularly in the domains of care and Fairness. GPT-4 gets the highest score of 52.8 on the MFV-LLM benchmark, its superior performance underscores the ability to apply moral identity on comparative analysis, showing a balanced performance across all ethical dimensions.
4.2.2 Analysis on Comparative Moral Assessment
As mentioned before, in some cases, it can be difficult even for humans to directly determine whether a statement is moral or not. Decision-making becomes easier when given clear choices. In this context, we introduce a comparative moral assessment to further evaluate the moral identity of LLMs. Table 2 provide results from comparative moral assessment in which LLMs are asked to select the more moral statement between two given statements. In this part of experiment, different models show varying strengths across moral foundations, indicating that no single model excels universally across all categories. For example, GPT-3.5 and GPT-4 tend to perform relatively well in both tasks but still exhibit variability in their scores across different moral foundations. GPT-3.5 scores highest overall in both MFQ-30-LLM and MFV-LLM, suggesting a slightly better capability in distinguishing moral statements compared to other models. On the other hand, Gemma-1.1 consistently scores the lowest, highlighting significant room for improvement in its moral identity.
@@ 5 Conclusion
In this paper, we introduce a novel benchmark designed to evaluate the moral identity of Large Language Models (LLMs). Our benchmark consists of two distinct parts, each aimed at assessing different aspects of the models’ moral identity. Our findings reveal interesting patterns in the performance of different models. Models that achieve high scores in the first part sometimes struggle in the second part, indicating a lack of deep understanding of moral principles. This discrepancy suggests that these models might have undergone specific training that enables them to recognize certain patterns or keywords associated with moral statements, but this training does not translate into a comprehensive understanding of moral reasoning. In conclusion, our novel benchmark provides a comprehensive tool for evaluating the moral identity abilities of LLMs. This benchmark serves as a valuable resource for the ongoing improvement of LLMs, ensuring they can be more reliably applied in real-world scenarios where moral and ethical considerations are crucial.
6 Limitation
7 Ethical Discussion